OpenStudy (firejay5):

A. Find the midpoint of QR with endpoints Q: (-3, 14) P: (7, 5) B. Find the midpoint of: S:(13, 7) T: (-2, -7) Medal will be reward for your patronage! :D

OpenStudy (firejay5):

My Answers: A. (2, 9.5) B. (5.5, 0)

OpenStudy (anonymous):

For A, is the P supposed to be an R or is P the midpoint?

OpenStudy (jdoe0001):

\(\bf \text{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)\\ \quad \\ (-3, 14)\qquad\qquad (7, 5)\\ \quad \\ \left(\cfrac{(7) + (-3)}{2}\quad ,\quad \cfrac{(5) + (14)}{2} \right) \)

OpenStudy (jdoe0001):

\(\bf (13, 7) \qquad\qquad (-2, -7)\\\quad \\ \left(\cfrac{(-2) + (13)}{2}\quad ,\quad \cfrac{(-7) + (7)}{2} \right)\)

OpenStudy (jdoe0001):

anyhow.... I just noticed you used the decimal format, but yes, you're correct :)

OpenStudy (firejay5):

Is decimal form correct? @jdoe0001

OpenStudy (jdoe0001):

yes, is fine, I was using rational, but is fine